OpenStudy (marziajackson$):

What is the area in square inches of a rectangle with a length of 6.3 feet, and a width of 2.5 feet? A. 189 in2 B. 15.75 in2 C. 141.75 in2 D. 2268 in2

Vocaloid (vocaloid):

Area = width*length 1 ft^2 = 12*12 inches^2 = 144 inches^2

OpenStudy (kash_thesmartguy):

\[6.3*2.5=15.75ft ^{2}\]It's B. Confirm it.

OpenStudy (johnweldon1993):

^If B were in ft^2 it would absolutely be correct...however we need to convert that to inches^2

OpenStudy (johnweldon1993):

Which as @Vocaloid has pointed out...can be done by multiplying your answer in ft^2 by 144in^2

OpenStudy (kash_thesmartguy):

You're wrong. This is just simple area. Just multiply the two sides together and there is your answer. BTW, 144 is not even an option.

OpenStudy (johnweldon1993):

Reread the question, the data and the answer choices again please...and I never said it was the answer...as we do not give away answers on this site

OpenStudy (johnweldon1993):

Now, @MarziaJackson$ we can make this easier by converting to inches right away...as we can see that the data presented was in ft...and we need inches for our answer... So \[\large 6.3ft \times \frac{12inches}{1ft} = 75.6 inches\] \[\large 2.5ft \times \frac{12inches}{1ft} = 30 inches\] So now...the area is simply \(\large 75.6in \times 30in = ?in^2\)

Vocaloid (vocaloid):

@Kash_TheSmartGuy: look at your units

Vocaloid (vocaloid):

the rectangle is 6.3 feet by 2.5 feet, which means that the area is 15.75 feet^2, like you said... HOWEVER, the answer choices are in inches^2 one square foot is 144 square inches, so how many square inches is 15.75 feet?